OpenStudy (anonymous):

evaluate without a calculator: logbase3(5) x logbase5(27) i changed it to 1/[logbase5(3)] x logbase5(27) then i got stuck D:

OpenStudy (paxpolaris):

it's multiplication not plus, huh?

OpenStudy (anonymous):

yup D:

OpenStudy (paxpolaris):

\[\log_3(5) \times \log_5 (27)\]\[={1 \over \log_5(3)} \times \log_5(27)\]

OpenStudy (paxpolaris):

27 = 3^3

OpenStudy (anonymous):

how did you change 1/logbase5(3) to 3^3?

jimthompson5910 (jim_thompson5910):

PaxPolaris is saying 27 is equal to 3^3 not 1/logbase5(3)

OpenStudy (anonymous):

oh yeah right oops so 1/logbase5(3) x logbase5(3^3) i don't see how i can find that D:

jimthompson5910 (jim_thompson5910):

what can you do with \[\Large \log_{5}(3^3)\]

jimthompson5910 (jim_thompson5910):

look at your notes for the log rules

OpenStudy (anonymous):

oh.. 1? so it becomes 1/logbase5(3) =logbase3(5) =?

jimthompson5910 (jim_thompson5910):

nope

jimthompson5910 (jim_thompson5910):

what can you do with that exponent?

OpenStudy (anonymous):

logbase5(3^3) is 1 isn't it

OpenStudy (paxpolaris):

\[\Large \log \left( a^b \right)=b \cdot \log \left( a \right)\]

jimthompson5910 (jim_thompson5910):

yes, so \[\Large \log_{5}(3^3) = ??\]

OpenStudy (anonymous):

oh okay 3logbase5(3) x 1/logbase5(3)

jimthompson5910 (jim_thompson5910):

and then you can say \[\Large \frac{3*\log_{5}(3)}{\log_{5}(3)}\]

jimthompson5910 (jim_thompson5910):

after you've multiplied the terms

OpenStudy (anonymous):

ah 3 thank you ^^

jimthompson5910 (jim_thompson5910):

yep